# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v5

@@ 5 Delegated Control and Collective Identity
The probability that the parts of a system embody a collective policy is higher if they embody weaker policies. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2025e].
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. A collective of cells remains cells regardless of whether I deem it to be an organ. The parts are still there. In terms of the stack, the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as w-maxing facilitates causal learning at the scale of an individual [bennett2023c], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
@@ 5.1 The Stack
From the enactive perspective, a human is something cells do, rather than are. An organ boils down to a collection of cells. The cells are an abstraction layer in which the collective policy of the organ is expressed, like software. However, software is just a state of hardware [bennett2024a]. Put another way, ‘organ’ is a label I ascribe to the the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs. ‘x86’ is a label I ascribe to a product of human behaviour. It is a static abstraction layer an engineer has embodied in silicon and then left alone; ‘fire and forget’. Humans act as the lower layer from which x86 emerges, but once it is embodied in silicon we disconnect it from our stack. We causally ‘intervene’ in the stack. To understand the effect of this intervention, I need to abstract away the lower levels. I must “abstract away” cells and focus on the consequences of their collective behaviour.
The cells are tasks. For tasks at higher levels of abstraction, capacity for w-maxing depends on lower levels. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ [^31] I can substitute $o$ for a declarative program $f$ that has the same extension, because that means $f$ and $o$ are true of the same states, having the same ‘truth conditions’ so to speak.. Hence I can transform policies to abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. I can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define ‘correctness’ at every level. At each successive level, ‘correct’ gets more specific. Because everything that happens in a higher level is in some sense correct according to lower levels, each higher level must more tightly constrain what is considered to be correct behaviour if that higher level is to exist in a meaningful sense.
@@
To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. I translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
In purely narrative terms, the implication of all this is that adaptation at higher levels depends on adaptation at lower levels. Below I state The Law of the Stack with a brief proof sketch; a full proof is provided in the Electronic Supplementary Materials.
@@
Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense:
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
@@
$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
In particular, increasing $\epsilon(\lambda_{i+1}(v_{i+1}))$ requires (at least) increasing $|E_{\pi_{i}}|$, i.e. weakening the policy at the lower layer.
@@
Utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$. Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$. By construction of the abstractor, $v_{i+1}$ is the image of $E_{\pi_{i}}$ under the map $o\mapsto\bigcap o$, hence $|v_{i+1}|\leq|E_{\pi_{i}}|$. Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|}$. ∎
@@ 5.2 Human Abstractions
Software might be agentic, but computer hardware is passive. Control is top-down. Once I have built an ISA, it exists in the environment independently of the human MCA from which it emerged. It is static. It cannot adapt, which is why there is demand for field programmable gate array (FPGA) [^32] An FPGA is like a CPU that can be reconfigured for specialised applications..
@@
If an abstraction layer is a window onto the world, then utility is how “good” the view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When I create an ISA I insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to what I can do with those abstractions I have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. This shows how adaptability at higher levels demands adaptability at lower levels. I am not suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [bennett2024b], because bounded systems can contain only finite information [bekenstein1981]. I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
@@ 6 Cancer-Like Failure Modes
I have explored some advantages of bio-mimetic distribution and delegation. Are there disadvantages? Yes. A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions. Some have argued cancer is a loss of collective identity [levin2021, davies2011]. In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity. When that happens, they independently pursue their own lower level goals. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. They escape from immune control mechanisms by becoming less differentiated, and closer to embryonic states.
@@
The set of correct policies shrinks as task difficulty increases, meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. If the policy at a lower level i1i-1 changes such that the task at level $i$ becomes ‘harder’333333i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in there being no correct policy at level $i$. If there are no longer any correct policies at level $i$, then the only way the system can continue to function is for part of the collective to ‘splinter’ or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind. Say I have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(\mathfrak{v}^{i-1}_{\mathfrak{a}},{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to ‘selective forgetting’ [bennett2023c] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, I would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that I obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of ‘correct’ policies to choose from.
@@
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [agosta2010], maintaining core functionality in as wide a variety of unanticipated circumstances as possible. The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Co-operation becomes impossible and the parts of a system start to compete. Conversely, if control of a system is delegated to the lowest levels, then adverse conditions are less likely to eliminate all correct policies. Put another way, there will be less incentive for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them. To summarise; a state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.
7 Hybrid Agency as Boundary-Condition Design

### v6

@@ 5 Delegated Control and Collective Identity
The probability that the parts of a system embody a collective policy is higher if they embody weaker policies. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [18].
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. A collective of cells remains cells regardless of whether I deem it to be an organ. The parts are still there. In terms of the stack, the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as w-maxing facilitates causal learning at the scale of an individual [14], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
@@ 5.1 The Stack
From the enactive perspective, a human is something cells do, rather than are. An organ boils down to a collection of cells. The cells are an abstraction layer in which the collective policy of the organ is expressed, like software. However, software is just a state of hardware [16]. Put another way, ‘organ’ is a label I ascribe to the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs. ‘x86’ is a label I ascribe to a product of human behaviour. It is a static abstraction layer an engineer has embodied in silicon and then left alone. ‘fire and forget’. Humans act as the lower layer from which x86 emerges, but once it is embodied in silicon we disconnect it from our stack. We causally ‘intervene’ in the stack. To understand the effect of this intervention, I need to abstract away the lower levels. I must “abstract away” cells and focus on the consequences of their collective behaviour.
The cells are tasks. For tasks at higher levels of abstraction, capacity for w-maxing depends on lower levels. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in\mathrm{Ext}(\pi)$ there exists $p\in P$ such that $T(o)=p$ (i.e. $p$ and $o$ are true of exactly the same states). Hence I can transform policies to abstraction layers, using a function $\mathfrak{f}$ that takes the extension $\mathrm{Ext}(\pi)$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{\,T(o):o\in\mathrm{Ext}(\pi)\,\}$. I can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define ‘correctness’ at every level. At each successive level, ‘correct’ gets more specific. Because everything that happens in a higher level is in some sense correct according to lower levels, each higher level must more tightly constrain what is considered to be correct behaviour if that higher level is to exist in a meaningful sense.
@@
To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. I translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $\mathrm{Ext}(\pi^{i})$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
In narrative terms, adaptation at higher levels depends on adaptation at lower levels. I now state The Law of the Stack and give a brief proof sketch. The ESM provides a full proof and additional lemmas.
@@
Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense.
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|\mathrm{Ext}(\pi_{i})|}.$
@@
$|\mathrm{Ext}(\pi_{i})|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
In particular, increasing $\epsilon(\lambda_{i+1}(v_{i+1}))$ requires (at least) increasing $|\mathrm{Ext}(\pi_{i})|$, i.e. weakening the policy at the lower layer.
Intuition. The bound says that higher-layer utility cannot grow without leaving enough slack in the realised policy one layer down. A narrow lower-layer policy induces a small next-layer vocabulary and this caps achievable outputs.
@@
Utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$. Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$. By construction of the abstractor, $v_{i+1}$ is the image of $\mathrm{Ext}(\pi_{i})$ under the map $o\mapsto T(o)$, hence $|v_{i+1}|\leq|\mathrm{Ext}(\pi_{i})|$. Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|\mathrm{Ext}(\pi_{i})|}$. ∎
Proof intuition. The proof sketch counts how many distinct higher-layer behaviours exist given the induced vocabulary at layer i+1i{+}1. The induced vocabulary is bounded by the number of completions left open by the realised policy at layer $i$.
5.2 Delegation and variational free energy
The Law of the Stack is a structural constraint. It says that degrees of freedom in higher layers is bounded by lower-layer slack. This aligns with the Free Energy Principle (FEP), which frames adaptive behaviour as variational free energy minimisation [39, 40, 71]. In the ESM (“Viability, lived history, and variational free energy”) I show that under a uniform viability prior, minimising free energy is equivalent to maximising the entropy (log-count) of viable future continuations. In Stack Theory terms, this is exactly w-maxing.
Combining that equivalence with Theorem 5.1 yields the corollary that reducing free energy at an abstract layer requires sufficient delegation (sufficiently weak realised policies) at lower layers. Intuitively, if lower layers are over-constrained, the induced higher-layer language is small, so even a maximum-entropy posterior cannot keep many viable futures open.
Corollary 5.1 (Delegation is necessary for lowering free energy).
Fix a planning horizon and let $\mu$ denote the viability task at layer i+1i{+}1. Write $O_{\mu}\subseteq L_{v_{i+1}}$ for the set of viable future continuations at that horizon. For any policy $\pi_{i+1}$ at layer i+1i{+}1, let $\Omega_{\pi_{i+1}}=\mathrm{Ext}(\pi_{i+1})\cap O_{\mu}$ be the set of viable futures compatible with $\pi_{i+1}$, and let $q_{\pi_{i+1}}$ be the uniform distribution on $\Omega_{\pi_{i+1}}$. Under the uniform viability prior assumed in the ESM, the corresponding (base-2) free energy satisfies
$F_{2}(q_{\pi_{i+1}})=\log_{2}|O_{\mu}|-\log_{2}|\Omega_{\pi_{i+1}}|.$
Moreover, $\Omega_{\pi_{i+1}}\subseteq L_{v_{i+1}}$, hence $|\Omega_{\pi_{i+1}}|\leq|L_{v_{i+1}}|\leq 2^{|\mathrm{Ext}(\pi_{i})|}$ as in the proof of Theorem 5.1, where $\pi_{i}$ is the realised policy at layer $i$ that induces $v_{i+1}$. Therefore,
$F_{2}(q_{\pi_{i+1}})\;\geq\;\log_{2}|O_{\mu}|-|\mathrm{Ext}(\pi_{i})|.$
In particular, free energy minimisation at layer i+1i{+}1 is bottlenecked by $|\mathrm{Ext}(\pi_{i})|$: reducing $F_{2}(q_{\pi_{i+1}})$ requires sufficient lower-layer slack, i.e. a weak enough realised policy at the lower layer.
Intuition. In active inference terms, free energy can only be reduced by keeping enough viable futures in play (high entropy). The Law of the Stack proves that if lower layers are frozen or over-constrained, there is an entropy bottleneck, so abstract-layer free energy has a floor.
@@ 5.3 Human Abstractions
Software might be agentic, but computer hardware is passive. Control is top-down. Once I have built an ISA, it exists in the environment independently of the human MCA from which it emerged. It is static. It cannot adapt, which is why there is demand for field programmable gate array (FPGA) (a reconfigurable processor substrate).
@@
If an abstraction layer is a window onto the world, then utility is how “good” the view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When I create an ISA I insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to what I can do with those abstractions I have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. This shows how adaptability at higher levels demands adaptability at lower levels. I am not suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [17], because bounded systems can contain only finite information [8]. I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
@@ 6 Cancer-Like Failure Modes
I have explored some advantages of bio-mimetic distribution and delegation. Are there disadvantages? Yes. A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions. Some have argued cancer is a loss of collective identity [57, 30]. In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity. When that happens, they independently pursue their own lower level goals. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. They escape from immune control mechanisms by becoming less differentiated, and closer to embryonic states.
@@
Formal characterisation of the cancer-analogue condition
To make the above precise, let $\alpha_{1},\dots,\alpha_{m}$ be $v$-tasks representing parts of a distributed system, with policy sets $\Pi_{\alpha_{j}}\subseteq L_{v}$, and let the collective task be $\omega$ with collective-policy set $\Pi_{\omega}:=\bigcap_{j=1}^{m}\Pi_{\alpha_{j}}$ (collective identity at $v$ requires $\Pi_{\omega}\neq\emptyset$).
Definition 6.1 (Restricted policies, over-constraint, and splinters).
To model top-down boundary conditions, fix restricted policy sets $\widehat{\Pi}_{\alpha_{j}}\subseteq\Pi_{\alpha_{j}}$ and define the restricted collective set $\widehat{\Pi}_{\omega}:=\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}$. We call the system over-constrained at $v$ (under the boundary conditions) if $\widehat{\Pi}_{\omega}=\emptyset$. A splinter is a non-empty index set $S\subseteq\{1,\dots,m\}$ such that $\bigcap_{j\in S}\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$.
Proposition 6.1 (Over-constraint implies loss of collective policy; any realised behaviour is splintered).
Assume at least one part remains locally viable, i.e. $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ for some $j$. If the system is over-constrained at $v$ (so $\widehat{\Pi}_{\omega}=\emptyset$), then there is no policy simultaneously feasible for all parts. Moreover, for any realised policy $\pi\in L_{v}$, the set of parts for which $\pi$ remains feasible,
$S(\pi):=\{\,j\in\{1,\dots,m\}:\pi\in\widehat{\Pi}_{\alpha_{j}}\,\},$
is a proper subset of $\{1,\dots,m\}$. In particular, any continued function can only proceed via a splinter (Definition 6.1).
Proof. If some $\pi$ were feasible for all parts then $\pi\in\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}=\widehat{\Pi}_{\omega}$, contradicting $\widehat{\Pi}_{\omega}=\emptyset$. Hence every realised policy fails at least one part’s constraints, so $S(\pi)\neq\{1,\dots,m\}$ for all $\pi\in L_{v}$. If $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ then $\{j\}$ is a splinter since $\bigcap_{k\in\{j\}}\widehat{\Pi}_{\alpha_{k}}=\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$. $\square$
As task difficulty increases (or top-down control tightens), the set of correct collective policies shrinks. A task $\alpha$ is easier than $\omega$ if $|\Pi_{\alpha}|>|\Pi_{\omega}|$. In Stack Theory terms, a cancer-like breakdown occurs when a policy change at layer i1i{-}1 induces a new collective task at layer $i$ with no correct policies, $\Pi_{\lambda^{i}(v^{i})}=\emptyset$, while components retain the capacity for independent action.
When $\Pi_{\lambda^{i}(v^{i})}=\emptyset$, the collective has no viable way to act as a coherent unit. Any continued function therefore requires dropping constraints, parent task splintering into ever narrower child tasks until those tasks are narrow enough to admit a policy. Part of the collective stops participating in the collective policy and reverts to a lower-level policy set with a much larger extension (more autonomy). This is a form of “selective forgetting” [14] applied to constraints rather than data.
@@
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [2], maintaining core functionality in as wide a variety of unanticipated circumstances as possible. The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Co-operation becomes impossible and the parts of a system start to compete. Conversely, if control of a system is delegated to the lowest levels, then adverse conditions are less likely to eliminate all correct policies. Put another way, there will be less incentive for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them. A state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.
7 Hybrid Creation as Weak Boundary-Condition Design